The Union Square Cinema charges $9 for an adult ticket and $4 for a child’s ticket. Last Saturday, the entire 1020 seat cinema was sold out. The total ticket receipts for the evening were $7680. How many of each type ticket were sold?
Answer by mananth(16949)

$9 ticket x numbers
$4 ticket y numbers

1 x + 1 y = 1020 .............1
9 x + 4 y = 7680 .............2

multiply (1)by -4
Multiply (2) by 1
-4 x -4 y = -4080
9 x + 4 y = 7680
Add the two equations
5 x = 3600
/ 5
x = 720
plug value of x in (1)
1 x + 1 y = 1020
720 + 1 y = 1020
1 y = 1020 -720
1 y = 300
y = 300
$ 9 tickets 720 numbers
$4 tickets 300 numbers
m.ananth@hotmail.ca